OpenStudy (anonymous):

RADICAL EQUATIONS -4√(x+9) = 20 (That is read at -4 times the square root of x+9 equals 20)

OpenStudy (anonymous):

16

OpenStudy (anonymous):

root(x + 9) = 20/-4 root(x + 9) = -5 squaring both sides x + 9 = 25 x = 25-9 x = 16

OpenStudy (anonymous):

Thats what I got when I did it but then I checked it and it didn't fit? When I checked it I got -5 ≠ 5.

OpenStudy (anonymous):

√(x+9)=-5 Square it x+9=25 so x=16

OpenStudy (anonymous):

@juicylovexx the checking works out root(x + 9) = 20/-4 root(x + 9) = -5 now put x = 16 and u get root(16 + 9) = -5 root(25) = -5 +/- 5 = -5 so -5 = -5 see 5² = 25 so root(25) = 5 also (-5)² = 25 s0 root(25) = -5

OpenStudy (anonymous):

Thank you! I get it now.

OpenStudy (anonymous):

That's not right harkirat. the square root is always positive. So this equation has no real solutions, I'm not sure about imaginary solutions.

OpenStudy (anonymous):

@Thomas9 when solving quadratic equations, we calculate +/- root(b² - 4ac) why do we take +/- values there????

OpenStudy (anonymous):

Yeap, no solutions exist

OpenStudy (anonymous):

That doesn't make the square root negative. http://www.wolframalpha.com/input/?i=solve+-4sqrt%28x%2B9%29%3D20

OpenStudy (anonymous):

so there is no solution?...